# Comparison results for Markov tree distributions

### v1

@@ 2.2 Stochastic orderings
Our comparison results are formulated in terms of integral stochastic orderings, which compare expectations of functions of two random vectors. Therefore, let $V=(V_{1},\ldots,V_{d})$ and $W=(W_{1},\ldots,W_{d})$ be $d$-variate random vectors defined on a probability space $(\Omega,\mathcal{A},P)$ which we assume to be non-atomic. For some class $\mathcal{F}$ of real-valued measurable functions $f\colon\mathbb{R}^{d}\to\mathbb{R}\,,$ the integral stochastic ordering
$V\prec_{\mathcal{F}}W\quad\text{is defined by}\quad\mathbb{E}f(V)\leq\mathbb{E}f(W)\quad\text{for all }f\in\mathcal{F},$
where the comparison of expectations is generally restricted to the subclass of functions in $\mathcal{F}$ such that the expectations exist. For $f\colon\mathbb{R}^{d}\to\mathbb{R}\,,$ denote by $\Delta_{i}^{\varepsilon}f(x):=f(x+\varepsilon e_{i})-f(x)$ the difference operator of length $\varepsilon>0$ applied to the $i$th component, where $e_{i}$ is the $i$th unit vector with respect to the standard base in $\mathbb{R}^{d}\,.$ Then $f$ is said to be supermodular respectively directionally convex, if $\Delta_{i}^{\varepsilon_{i}}\Delta_{j}^{\varepsilon_{j}}f(x)\geq 0$ for all $x\in\mathbb{R}^{d}$ and for all $1\leq i<j\leq d$ and $1\leq i\leq j\leq d\,,$ respectively. Further, $f$ is said to be $\Delta$-monotone respectively $\Delta$-antitone, if $\Delta^{\varepsilon_{1}}_{i_{1}}\cdots\Delta^{\varepsilon_{k}}_{i_{k}}f(x)\geq 0$ and $(-1)^{k}\Delta^{\varepsilon_{1}}_{i_{1}}\cdots\Delta^{\varepsilon_{k}}_{i_{k}}f(x)\geq 0\,,$ respectively, for all $k\in\{1,\ldots,d\}$, $\varepsilon_{1},\ldots,\varepsilon_{k}>\nolinebreak 0$, $\{i_{1},\ldots,i_{k}\}\subseteq\{1,\ldots,d\}$ and $x\in\mathbb{R}^{d}$. Note that $\Delta$-monotone functions and $\Delta$-antitone functions are supermodular. Further, directionally convex functions are the functions that are supermodular and componentwise convex. If $f$ is sufficiently smooth, then $f$ is supermodular if and only if $\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f(x)\geq 0$ for all $1\leq i<j\leq d$ and for all $x\,.$ Similar properties hold true for sufficiently smooth directionally convex, $\Delta$-monotone and $\Delta$-antitone functions [87]. For several examples of such functions, see Table 1.
@@
Denote by $\mathcal{F}_{\uparrow}\,,$ $\mathcal{F}_{sm}\,,$ $\mathcal{F}_{dcx}\,,$ the class of componentwise increasing, supermodular and directionally convex functions, respectively, by $\mathcal{F}_{\uparrow}\,,$ $\mathcal{F}_{cx}\,,$ and $\mathcal{F}_{icx}:=\mathcal{F}_{\uparrow}\cap\mathcal{F}_{cx}\,$ the class of increasing, convex, and increasing convex functions on $\mathbb{R}$, respectively. Let $F_{V}$ and $\overline{F}_{V}$ be the distribution function and survival function associated with a random vector $V\,.$ We make use of the following integral stochastic orderings.
@@
    Let $V=(V_{1},\ldots,V_{d})$ and $W=(W_{1},\ldots,W_{d})$ be $d$-variate random vectors. Then $V$ is said to be smaller than $W$ with respect to
@@
      the lower orthant order, written $V\leq_{lo}W$, if $F_{V}(x)\leq F_{W}(x)$ for all $x\in\mathbb{R}^{d}\,,$
@@
      the upper orthant order, written $V\leq_{uo}W$, if $\overline{F}_{V}(x)\leq\overline{F}_{W}(x)$ for all $x\in\mathbb{R}^{d}\,,$
@@
      the concordance order, written $V\leq_{c}W$, if $V\leq_{lo}W$ and $V\leq_{uo}W\,,$
@@
    5. (v)
      the directionally convex order, written $V\leq_{dcx}W$, if $V\prec_{\mathcal{F}_{dcx}}W$,
    6. (vi)
      the stochastic order, written $V\leq_{st}W$, if $V\prec_{\mathcal{F}_{\uparrow}}W$.
@@
  3. (c)
    Let $S$ and $T$ be real-valued random variables. Then $S$ is said to be smaller than $T$ with respect to
    1. (vii)
      the convex order, written $S\leq_{cx}T$, if $S\prec_{\mathcal{F}_{cx}}T$,
    2. (viiii)
      the increasing convex order, written $S\leq_{icx}T$, if $S\prec_{\mathcal{F}_{icx}}T$.
@@
Note that the comparison of stochastic processes in Definition 2.7(b) is defined through the comparison of the finite-dimensional marginal distributions, which corresponds to the notion of strong comparison of stochastic processes in [87, Definition 5.1.2]. Further, the lower orthant order and the upper orthant order are also integral stochastic orderings which are generated by the class of $\Delta$-antitone and $\Delta$-monotone functions, i.e.,
$\begin{split}V\leq_{lo}W\quad&\Longleftrightarrow\quad\mathbb{E}f(V)\leq\mathbb{E}f(W)\quad\text{for all }f\in\mathcal{F}_{\Delta}^{-}\,,\\ V\leq_{uo}W\quad&\Longleftrightarrow\quad\mathbb{E}f(V)\leq\mathbb{E}f(W)\quad\text{for all }f\in\mathcal{F}_{\Delta}\,,\end{split}$ (7)
see [95, 101]. Some basic relations between the above considered integral stochastic orderings are
[…]
@@
As a direct consequence of the definition, the concordance order requires that $V$ and $W$ have the same univariate marginal distributions, i.e., $V\leq_{c}W$ implies $V_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{i}$ for all $i\in\{1,\ldots,d\}\,.$ Due to (8), also the supermodular ordering requires equal marginal distributions and, thus, both $\leq_{c}$ and $\leq_{sm}$ are pure dependence orders. In particular, for bivariate random vectors $V$ and $W$ with the same univariate marginal distributions, the lower orthant, upper orthant, concordance, supermodular and directionally convex order are equivalent, i.e., if $d=2\,,$ then
$V\leq_{lo}W$$\Longleftrightarrow$ $V\leq_{uo}W$$\Longleftrightarrow$ $V\leq_{c}W$$\Longleftrightarrow$ $V\leq_{sm}W$$\Longleftrightarrow$ $V\leq_{dcx}W$ (9)
whenever $V_{1}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{1}$ and $V_{2}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{2}\,,$ see [84, Theorem 2.5]. Further, $V\leq_{c}W$ and thus $V\leq_{sm}W$ implies $\mathrm{Cor}(V_{i},V_{j})\leq\mathrm{Cor}(W_{i},W_{j})$ for all $i\neq j\,,$ where $\mathrm{Cor}$ denotes the correlation in the sense of Pearson, Spearman or Kendall, whenever defined, see [101, Remark 6.3]. Since, for $d\geq 3\,,$ $V\leq_{c}W$ does not imply $\sum_{i}V_{i}\leq_{cx}\sum_{i}W_{i}$ and thus also not $V\leq_{sm}W\,,$ see [84, Theorem 2.6], we focus on comparison results with respect to the stronger concept of supermodular order. As a consequence, we obtain inequalities for various classes of functionals relevant to many applications such as
$V\leq_{sm}W\quad\text{or}\quad V\leq_{dcx}W$$\Longrightarrow$ $h(V)\leq_{icx}h(W)$$\Longrightarrow$ $\Psi\Big{(}h(V)\Big{)}\leq\Psi\Big{(}h(W)\Big{)},$ (10)
where $h$ is a componentwise increasing or decreasing supermodular function and $\Psi$ is a convex, law-invariant risk measure on a proper space such as the space of integrable or the space of bounded random variables, see [101, Corollary 6.16], [47, Chapter 4] and [15, 28, 64, 100]. If $h(V)=\sum_{i=1}^{d}V_{i}$, then $V\leq_{sm}W$ or $V\leq_{dcx}W$ implies $\sum_{i=1}^{d}V_{i}\leq_{cx}\sum_{i=1}^{d}W_{i}\,,$ i.e., the component sums are then ordered with respect to the convex order. Hence, due to (10), supermodular or directionally convex comparison results yield, in particular, various comparison results for risk functionals such as the average-value-at-risk of the aggregated risk vector, which may stand for a portfolio risk in finance or for the risk of total damages in insurance. The supermodular order also has the important property that it is invariant under increasing transformations of the components, i.e., for all increasing functions $k_{1},\ldots,k_{d}\colon\mathbb{R}\to\mathbb{R}\,,$ one has
$(V_{1},\ldots,V_{d})\leq_{sm}(W_{1},\ldots,W_{d})$$\Longrightarrow$ $(k_{1}(V_{1}),\ldots,k_{d}(V_{d}))\leq_{sm}(k_{1}(W_{1}),\ldots,k_{d}(W_{d})).$ (11)
Further, in the pure marginal model, the comonotonic random vector $V^{c}=(V_{1}^{c},\ldots,V_{d}^{c})\,,$ $V_{i}^{c}:=F_{V_{i}}^{-1}(U)$ for all $i$ with $U$ uniformly distributed on $(0,1)\,,$ describes the worst case distribution in supermodular order, i.e.,
$V\leq_{sm}V^{c}\,,$ (12)
see, e.g., [87, Theorem 3.9.14]. For an overview of stochastic orderings, we refer to [87, 101, 104].
2.3 Positive dependence concepts
@@
For modeling positive dependencies, we make use of several positive dependence concepts. To this end, for a $d$-variate random vector $V=(V_{1},\dots,V_{d})$, denote by $V^{\perp}:=(V_{1}^{\perp},\ldots,V_{d}^{\perp})$ an independent random vector with the same marginal distributions as $V\,,$ i.e., $V_{1}^{\perp},\ldots,V_{d}^{\perp}$ are independent and $V_{i}^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}V_{i}$ for all $i\in\{1,\dots,d\}$.
@@
    positive lower orthant dependent (PLOD) if $V^{\perp}\leq_{lo}V\,,$
@@
    positive upper orthant dependent (PLOD) if $V^{\perp}\leq_{uo}V,$
@@
    conditionally increasing (CI) if
    $V_{i}\uparrow_{st}(V_{j}\,,\leavevmode\nobreak\ j\in J)$ (13)
  5. (v)
  6. (vi)
    multivariate totally positive of order $2$ ($\mathrm{\text{MTP}_{2}}$) if $V$ is absolutely continuous with Lebesgue-density $f$ such that $\log(f)$ is supermodular.
The above concepts of positive dependence are defined similarly for probability distributions and distribution functions, and they are invariant under increasing transformations, see [87, Theorem 3.10.19]. Hence, positive dependence is a copula-based property.
@@
A simple criterion which is equivalent to condition (13) is that the conditional survival functions are pointwise increasing, i.e., for $J=(j_{1},\ldots,j_{m})$ and for all $y\in\mathbb{R}\,,$ the conditional probability $P(V_{i}>y\mid V_{j_{1}}=x_{1},\ldots,V_{j_{m}}=x_{m})$ is increasing in $x_{1},\ldots,x_{m}$ outside a $V_{J}$-null set, see, e.g., [87, Example 2.5.2]. Hence, a bivariate random vector $(V_{1},V_{2})$ is CIS (i.e., $V_{2}\uparrow_{st}V_{1}$) if and only if, for all $x_{2}\in\mathbb{R}\,,$ $P(V_{2}>x_{2}\mid V_{1}=x_{1})$ is increasing in $x_{1}$ outside a $V_{1}$-null set which may depend on $x_{2}\,.$ A sufficient (and in the case of continuous marginal distributions also necessary) criterion for the latter property is that the underlying copula is concave in its first component, which follows from the Sklar-type representation of conditional distribution functions in [6, Theorem 2.2].
The above defined positive dependence concepts are related by
[…]
where the first three implications are strict for $d\geq 2$ and the last implication is strict only for $d\geq 3$, see [87, page 146] for an overview of these concepts. Note that PLOD or PUOD imply pairwise non-negative correlations in the sense of Pearson, Kendall, and Spearman, whenever these coefficients are defined, see, e.g., [101, Remark 6.3].
3 Ordering results for Markov tree distributions
In this section, we provide integral inequalities $\mathbb{E}f(X)\leq\mathbb{E}f(Y)$ for Markov tree distributed sequences $X\sim\mathcal{M}(F,T,B)$ and $Y\sim\mathcal{M}(G,T,C)$ with respect to the marginal specifications $F=(F_{n})_{n\in N}$, $G=(G_{n})_{n\in N}$ and the bivariate dependence specifications $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E},$ respectively. First, we give a variant of Theorem 1.3 formulated in terms of bivariate tree specifications. Then, we establish general criteria on the marginal specifications $F$, $G$ and the bivariate copula families $B$, $C$ leading to comparison results with respect to the lower orthant, upper orthant, and directionally convex order. Care is required when comparing Markov tree distributions with different and discontinuous marginal distributions since conditional independence is not a pure dependence property, see Example A.2. All proofs of this section are deferred to Appendix D.
3.1 Inequalities for classes of supermodular functions
Due to Proposition 2.5, every Markov tree distribution has a representation by a bivariate tree specification, i.e., by a family of univariate distribution functions and a family of bivariate copulas specifying the nodes and the edges of the underlying directed tree. Vice versa, every bivariate tree specification has a unique Markov tree realization. In this light, Theorem 1.3 can be translated into the notation of bivariate tree specifications, which often proves useful in practice. To this end, assume that $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ follow Markov tree distributions with respect to a tree $T=(N,E)\,.$ Since the supermodular order is a pure dependence order and implies that the respective marginal distributions coincide, we obtain from condition (iii) of Theorem 1.3, i.e., $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E$, that
$F_{X_{n}}=F_{Y_{n}}=:F_{n}\quad\text{for all }n\in N\,.$
Further, due to identical marginal distributions, there exist by (9) bivariate copulas for $(X_{i},X_{j})$ and $(Y_{i},Y_{j})$ which are pointwise ordered, i.e.,
$C_{X_{i},X_{j}}\leq_{lo}C_{Y_{i},Y_{j}}\quad\text{for all }e=(i,j)\in E\,,$
Condition (i) of Theorem 1.3, i.e., $X_{j}\uparrow_{st}X_{i}\,,$ $(i,j)\in E$, means that the bivariate random vector $(X_{i},X_{j})$ is conditionally increasing in sequence (CIS), see Remark 2.9. Using that the CIS property is invariant under increasing transformations, it follows that there exists a bivariate copula $B_{e}$ such that
$B_{e}\text{ is CIS}\quad\text{and}\quad B_{e}=C_{X_{i},X_{j}}\text{ on }\mathsf{Ran}(F_{i})\times\mathsf{Ran}(F_{j})\quad\text{for }e=(i,j)\in E\,.$
Recall that a bivariate copula is CIS if and only if it is concave in the first component. Denote by $D^{\top}$ the transpose of a bivariate copula $D\,,$ i.e., $D^{\top}(u,v):=D(v,u)$ for all $(u,v)\in[0,1]^{2}\,.$ Then, condition (ii) of Theorem 1.3, i.e., $Y_{i}\uparrow_{st}Y_{j}$ (and $Y_{j}\uparrow_{st}Y_{i}$), $(i,j)\in E$, means that there exists a bivariate copula $C_{e}$ such that
$C_{e}^{\top}\text{ is {CIS} ($C_{e}$ is {CI})}\quad\text{and}\quad C_{e}=C_{Y_{i},Y_{j}}\text{ on }\mathsf{Ran}(F_{i})\times\mathsf{Ran}(F_{j})\,.$
Finally, we have decomposed $X$ and $Y$ into the bivariate tree specification with marginal distribution functions $F=(F_{n})_{n\in N}$ and pointwise ordered CIS (or CI) copulas $B=(B_{e})_{e\in E}$ and $C=(C_{e})_{e\in E}$. Since members of many well-known bivariate copula families are CI, see [5], which implies CIS, we formulate the following variant of Theorem 1.3 under slightly stronger assumptions on the bivariate specifications of $X$ and $Y\,.$
Theorem 3.1 (Supermodular ordering based on bivariate tree specifications)
Let $X\sim\mathcal{M}(F,T,B)$ and $Y\sim\mathcal{M}(F,T,C)$ be Markov tree distributions. Assume for all $e\in E$ that

### v3

@@ 2.2 Stochastic orderings and positive dependence concepts
Our comparison results are formulated in terms of integral stochastic orderings, which compare expectations of functions of two random vectors. To this end, consider for a measurable weight function $b\colon\mathbb{R}^{d}\to[1,\infty)$ (e.g., $b(x)=\max\{1,\lVert x\rVert^{p}\}$ for some $p\in\mathbb{N}_{0}$) the class $\mathcal{R}_{b}:=\{V=(V_{1},\ldots,V_{d})\mid\mathbb{E}b(V)<\infty\}$ of $b$-integrable random vectors on a probability space $(\Omega,\mathcal{A},P).$ Then, for some class $\mathcal{F}\subseteq\{f\colon\mathbb{R}^{d}\to\mathbb{R}\mid f\text{ measurable},\sup_{x\in\mathbb{R}^{d}}|f(x)|/b(x)<\infty\}$ of $b$-bounded functions and for $V,W\in\mathcal{R}_{b},$ define the integral stochastic order
$V\prec_{\mathcal{F}}W\quad\text{if}\quad\mathbb{E}f(V)\leq\mathbb{E}f(W)\quad\text{for all }f\in\mathcal{F};$
see [56]. Recall that a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is supermodular if $f(x)+f(y)\leq f(x\wedge y)+f(x\vee y)$ for all $x,y\in\mathbb{R}^{n}.$ The function $f$ is said to be increasing/decreasing supermodular if it is supermodular and componentwise increasing/decreasing. If $f$ is supermodular and componentwise convex, it is directionally convex. For several examples of such functions, see Table 1. If $f$ is sufficiently smooth, then $f$ is supermodular if and only if $\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f(x)\geq 0$ for all $1\leq i<j\leq d$ and for all $x.$ Similar properties hold true for sufficiently smooth increasing/decreasing supermodular or directionally convex functions.
@@
Denote by $\mathcal{F}_{\uparrow},$ $\mathcal{F}_{sm},$ $\mathcal{F}_{ism},$ $\mathcal{F}_{dsm},$ and $\mathcal{F}_{dcx},$ the class of componentwise increasing, supermodular, increasing supermodular, decreasing supermodular, and directionally convex $b$-bounded functions, respectively. We will use the following integral stochastic orderings.
@@
    Let $V=(V_{1},\ldots,V_{d}),W=(W_{1},\ldots,W_{d})\in\mathcal{R}_{b}$. Then $V$ is said to be smaller than $W$ with respect to
@@
      the stochastic order, written $V\leq_{st}W$, if $V\prec_{\mathcal{F}_{\uparrow}}W$,
@@
      the increasing (decreasing) supermodular order, written $V\leq_{ism}W$ ($V\leq_{dsm}W$), if $V\prec_{\mathcal{F}_{ism}}W$ ($V\prec_{\mathcal{F}_{dsm}}W$),
@@
      the directionally convex order, written $V\leq_{dcx}W$, if $V\prec_{\mathcal{F}_{dcx}}W$.
@@
For $d=1,$ the directionally convex order reduces to the well known convex order $\leq_{cx},$ and the increasing convex order $\leq_{icx}$ is defined for the class of functions that are increasing and convex. Note that the comparison of stochastic processes in part (b) is defined through the comparison of their finite-dimensional marginal distributions. This corresponds to the concept of strong comparison of stochastic processes in [61, Definition 5.1.2]. We denote by $V\leq_{lo}W$ and $V\leq_{uo}W$ the lower and upper orthant order, defined by $F_{V}(x)\leq F_{W}(x)$ and $\overline{F}_{V}(x)\leq\overline{F}_{W}(x)$ for all $x\in\mathbb{R}^{d},$ respectively. Some basic relations between the above considered orderings are
$V\leq_{dcx}W\penalty 10000\ \text{and}\penalty 10000\ V\leq_{ism}W\quad\Longleftarrow\quad V\leq_{sm}W\penalty 10000\ \penalty 10000\ \Longrightarrow\quad V\leq_{lo}W\penalty 10000\ \text{and}\penalty 10000\ V\leq_{uo}W,$ (4)
$V\leq_{ism}W\quad\Longrightarrow\quad\max_{n}\{V_{n}\}\geq_{st}\max_{n}\{W_{n}\}\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \sum_{n}V_{n}\leq_{icx}\sum_{n}W_{n}.$ (5)
@@
Further, $V\leq_{dcx}W$ implies $\sum_{n}V_{n}\leq_{cx}\sum_{n}W_{n}$. If $\mathbb{E}[V_{n}]=\mathbb{E}[W_{n}]$ for all $n,$ the increasing supermodular order reduces to the supermodular order and the increasing convex order in (5) reduces to the convex order. As a direct consequence of (4), the supermodular order is a pure dependence order, i.e., $V\leq_{sm}W$ implies $V_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{i}$ for all $i\in\{1,\ldots,d\},$ where $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$ denotes equality in distribution. For bivariate random vectors with identical marginal distributions, the orderings in (4) are equivalent, i.e., if $d=2,$ then
$V\leq_{lo}W$ $\Longleftrightarrow$ $V\leq_{uo}W$ $\Longleftrightarrow$ $V\leq_{sm}W$ $\Longleftrightarrow$ $V\leq_{dcx}W$ (6)
whenever $V_{1}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{1}$ and $V_{2}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}W_{2};$ see [58, Theorem 2.5]. Hence, for identical marginals, a verification of the ordering condition (iii) in Theorem 1.3 reduces to a pointwise comparison of the associated bivariate distribution functions or copulas.
@@
For modeling positive dependencies, we make use of several positive dependence concepts. To this end, we denote by $V^{\perp}:=(V_{1}^{\perp},\ldots,V_{d}^{\perp})$ an independent random vector with the same marginal distributions as $V=(V_{1},\dots,V_{d})$, i.e., $V_{1}^{\perp},\ldots,V_{d}^{\perp}$ are independent and $V_{i}^{\perp}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}V_{i}$ for all $i\in\{1,\dots,d\}$.
@@
    positive supermodular dependent (PSMD) if $V^{\perp}\leq_{sm}V,$
@@
    conditionally increasing (CI) if
    $V_{i}\uparrow_{st}(V_{j},\penalty 10000\ j\in J)$ (7)
@@
    multivariate totally positive of order 22 ($\mathrm{\text{MTP}_{2}}$) if $V$ is absolutely continuous with Lebesgue-density $f$ such that $\log(f)$ is supermodular.
The above concepts of positive dependence are defined similarly for probability distributions and distribution functions, and remain invariant under (strictly) increasing transformations; see [61, Theorem 3.10.19] and [17]. The concepts are related by
$\mathrm{\text{MTP}_{2}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Longrightarrow\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{CI}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Longrightarrow\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{CIS}\penalty 10000\ \penalty 10000\ \penalty 10000\ \Longrightarrow\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{PSMD},$ (8)
where the implications are strict for dimension $\geq 2.$
3 Flexibility in the marginal distributions
In Theorem 1.3 and Corollary 1.4, we have compared Markov tree distributions with respect to the supermodular order, which requires identical marginal distributions. However, if the marginals of $X$ and $Y$ differ or are only partially known, some flexibility of our results with respect to the marginals is desirable. First, we study flexibility of the marginal distributions in stochastic order, which yields a comparison result with respect to the increasing/decreasing supermodular order. Then we allow flexibility of the marginal distributions in convex order and obtain a comparison result with respect to the directionally convex order. Since conditional independence is not solely a copula-based property, caution is required when comparing Markov tree distributions with different and discontinuous marginal distribution functions; see Example 6.4.
3.1 Inequalities for increasing/decreasing supermodular functions
In the following result, we additionally compare the marginal distributions in stochastic order. We assume that the ranges of the marginal distribution functions are identical except on Lebesgue null sets. The closure of a set $A\subseteq[0,1]$ is denoted by $\overline{A}.$
Theorem 3.1 (Comparison of marginals in stochastic order)
Assume that $X=(X_{n})_{n\in N}\sim\mathcal{M}(F,T,B)$ and $Y=(Y_{n})_{n\in N}\sim\mathcal{M}(G,T,C)$ satisfy the positive dependence conditions (i) and (ii) of Theorem 1.3 and assume that $\overline{\mathsf{Ran}(F_{n})}=\overline{\mathsf{Ran}(G_{n})}$ for all $n\in N.$ Then the following statements hold true:
  1. (i)
    $B_{e}\leq_{lo}C_{e}$ for all $e\in E$ and $F_{n}\leq_{st}G_{n}$ for all $n\in N$ implies $X\leq_{ism}Y,$
  2. (ii)
    $B_{e}\leq_{lo}C_{e}$ for all $e\in E$ and $F_{n}\geq_{st}G_{n}$ for all $n\in N$ implies $X\leq_{dsm}Y.$
@@ 3.1 Inequalities for increasing/decreasing supermodular functions
  1. (a)
    As a consequence of Theorem 3.1, we obtain distributional robustness for various functionals that are consistent with the increasing resp. decreasing supermodular order. In particular, we obtain distributional robustness of extreme order statistics and sums of random variables with respect to the stochastic order and the increasing convex order as in (5). Note that for bivariate copulas, $B_{e}\leq_{lo}C_{e}$, $B_{e}\leq_{uo}C_{e}$, and $B_{e}\leq_{sm}C_{e}$ are equivalent due to (6).
  2. (b)
    Due to Example 6.4, the assumption $\overline{\mathsf{Ran}(F_{n})}=\overline{\mathsf{Ran}(G_{n})}$ cannot be omitted because the univariate marginal distributions generally affect the dependence structure under the Markov property. For the proof of Theorem 3.1, we use that the copula of a Markov tree distribution depends only on the ranges of the marginal distribution functions.
  3. (c)
    The assumption $\overline{\mathsf{Ran}(F_{n})}=\overline{\mathsf{Ran}(G_{n})}$ is, in particular, satisfied when $F_{n}$ and $G_{n}$ are continuous. In Example 5.3, we use the fact that the random variables $\xi_{i}^{+}:=\max\{\xi_{i},0\},$ for $\xi_{i}\sim N(0,\sigma_{i}),$ $i=1,2,$ are comparable in stochastic order and satisfy $\mathsf{Ran}(F_{\xi_{1}^{+}})=\mathsf{Ran}(F_{\xi_{2}^{+}})=\{0\}\cup[1/2,1).$ More precisely, if $\sigma_{1}\leq\sigma_{2},$ then $\xi_{1}^{+}\leq_{st}\xi_{2}^{+}.$
function $f(x_{1},\ldots,x_{d})$ Properties
$\mathds{1}_{\{x\geq t\}}$ increasing supermodular
$\mathds{1}_{\{x\leq t\}}$ decreasing supermodular
$P(\xi_{1}\leq x_{1},\ldots,\xi_{d}\leq x_{d})$ increasing supermodular
$P(\xi_{1}\geq x_{1},\ldots,\xi_{d}\geq x_{d})$ decreasing supermodular
$\min\{x_{1},\ldots,x_{d}\}$ increasing supermodular
$\max\{x_{1},\ldots,x_{d}\}$ increasing supermodular, directionally convex
$\mathds{1}_{\{\min\{x_{1},\ldots,x_{d}\}\geq K\}}$ increasing supermodular
$\mathds{1}_{\{\max\{x_{1},\ldots,x_{d}\}\leq K\}}$ decreasing supermodular
$(K-\max\{x_{1},\ldots,x_{d}\})_{+}$ decreasing supermodular
$(\max\{x_{1},\ldots,x_{d}\}-K)_{+}$ increasing supermodular, directionally convex
$(K-\min\{x_{1},\ldots,x_{d}\})_{+}$ decreasing supermodular, directionally convex
$(\min\{x_{1},\ldots,x_{d}\}-K)_{+}$ increasing supermodular
$\varphi(\sum_{n=1}^{d}\alpha_{n}x_{n})$, $\varphi$ convex supermodular, directionally convex
$\varphi(\sum_{n=1}^{d}\alpha_{n}x_{n})$, $\varphi$ increasing convex increasing supermodular, directionally convex
$\varphi(\sum_{n=1}^{d}g_{n}(x_{n}))$, $\varphi$ convex supermodular
$\varphi(\sum_{n=1}^{d}g_{n}(x_{n}))$, $\varphi$ increasing convex increasing supermodular
Table 1: Important examples and classifications of functions implying comparison results for $\mathbb{E}f(X)$ with respect to the integral stochastic orders considered in this paper, where $(\xi_{1},\ldots,\xi_{d})$ is a random vector (independent of $X$) on a probability space $(\Omega,\mathcal{A},P)$ and where $t\in\mathbb{R}^{d},$ $K\in\mathbb{R},$ $(y)_{+}:=\max\{y,0\},$ $\alpha_{n}\geq 0,$ and $g_{n}\colon\mathbb{R}\to\mathbb{R}$ increasing. Since the integral stochastic orders satisfy various invariance properties, the above examples also apply to the respective transformations of the functions, for example to componentwise increasing/decreasing transformations in the case of the supermodular order; see [13].
3.2 Inequalities for directionally convex functions
When the marginals are in convex order, a comparison with respect to the directionally convex order can be achieved if the underlying multivariate copula is CI. This is the content of the following well known result, which is not restricted to Markov tree distributions.
Lemma 3.3 (Common CI copula; [59, Theorem 4.5])
Let $U=(U_{0},\ldots,U_{d})$ and $V=(V_{0},\ldots,V_{d})$ be random vectors having the same $(d+1)$-dimensional copula $C=C_{U}=C_{V}.$ If $C$ is CI, then $U_{i}\leq_{cx}V_{i}$ for all $i\in\{0,\ldots,d\}$ implies $U\leq_{dcx}V.$
As we demonstrate in Example 6.5, bivariate CI specifications generally do not result in a Markov tree distribution with a CI copula. A sufficient condition for the underlying copula to be CI is provided by the following special case of [32, Proposition 7.1].
Lemma 3.4 ($\mathrm{\text{MTP}_{2}}$ specifications)
Let $V=(V_{0},\ldots,V_{d})\sim\mathcal{M}(G,T,C)$ be a Markov tree distributed random vector such that $(V_{i},V_{j})$ is $\mathrm{\text{MTP}_{2}}$ for all $e=(i,j)\in E.$ Then $V$ is $\mathrm{\text{MTP}_{2}}.$
Combining the above lemmas and Theorem 1.3, we establish the following $\leq_{dcx}$-comparison result for Markov tree distributions, which allows a comparison of the marginal distributions in convex order. Note that the bivariate copula specifications of $Y$ are now assumed to satisfy the stronger positive dependence concept $\mathrm{\text{MTP}_{2}}.$
Theorem 3.5 (Comparison of marginals in convex order)
For a tree $T=(N,E),$ let $X=(X_{n})_{n\in N}\sim\mathcal{M}(F,T,B)$ and $Y=(Y_{n})_{n\in N}\sim\mathcal{M}(G,T,C)$ be Markov tree distributed sequences of random variables. Assume for all $n\in N$ that the marginal distribution functions $F_{n}$ and $G_{n}$ are continuous. If for all $e=(i,j)\in E,$